OpenStudy (anonymous):

WHICH IS AN EQUATION IN POINT-SLOPE FORM OF THE LINE THAT PASSES THROUGH THE POINTS (7,5) AND (-4-,-1) A.Y+1=6/7(X+4) B.Y+1=6/11(X+4) C.Y+4=11/6(X+1) D.Y-1=6/11(X+4)

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

point slope form is y-y1=m(x-x1) m=(y2-y1)/(x2-x1) (5-(-1))/(7-(-4)) 6/11 y+1=6/11(x+4)